OpenStudy (anonymous):

Find the equation of the line containing the points (-9, -2) and (3,4)

hero (hero):

y = mx + b -2 = -9m + b 4 = 3m + b 9m - 2 = b 4 - 3m = b b = b 9m - 2 = 4 - 3m 9m + 3m = 4 + 2 12m = 6 m = 6/12 m = 1/2 b = 9(1/2) - 2 b = 4 1/2 - 2 b = 2 1/2 b = 5/2 y = (1/2)x + 5/2 y = (x + 5)/2